From GlueXWiki
Jump to: navigation, search

June 9, 2011 FDC meeting


  1. Production (Dave)
    • First chamber assembly: HV problem
    • Spacer ring: gas leakage
    • Second set wire frame and cathodes
    • Other
  2. Engineering (Bill)
    • Vacuum chamber
    • Ultrasonic cleaning
    • Rigid-flex gluing tool
    • Other
  3. Electronics (Fernando, Chris)
    • PCB schematics update
    • Other
  4. Chamber testing
    • Wire Position Measurements
  5. Other


Participants: Eugene, Bill, Dave, Chris, Caleb, Beni, Simon, and Lubomir.


- End of last week we assembled the first production chamber, run gas over the weekend, on Monday applied HV and was tripping (1uA threshold) at very low voltages: ~20-30V for sense wires and somewhat higher (~30V, one sector was at -500V for short time) for field wires. Then we found ~1-10MOhm between HV and ground and practically all the pads on the PCBs, even between opposite PCBs on the wire frame that have no electrical connection. The conclusion was: the wire frame got wet (most likely the Rohacell) when it was cleaned in the ultrasonic bath. We tried to dry it with a lamp and then with a heat gun; didn't work. Then today we put it a big vacuum chamber at the Test Lab to take the moisture out. The person servicing the vacuum chamber mentioned that it went down to 10^-4 torrs for ~2hours, much slower than usual (~20min) which indicates there was a lot of moisture there. We will keep it there by tomorrow morning when we expect 10^-7 torrs. Bill: such under pressure will not damage the Rohacell.

- The good news is that the gas leakage of the first package was very low (undetectable) except at one place: one of the tube on the spacer ring. Mark Stevens was yesterday at Blue Crab explaining the procedure for gluing the gas tubing on a new spacer ring; the first one will be fixed too.

- We are working now on the second chamber set. The second wire frame was strung, glued, positions measured, then the wires were soldered. (After the meeting also the tensions were measured). Tomorrow the wires will be cut and the PCBs populated. Cathodes #3 and #4 are type 3 (as #1 and #2) and ready mechanically. After testing the modified tool we will glue the rigid-flexes.

- Starting the beginning of this week, the AC of the bay area was set to higher temperature (85 F) only over the night. Ron Bartek monitors the work of the clean room AC (settings unchanged). Eugene was concerned also about the procedures and materials that are outside of the clean room. So far the humidity there was not so high (45%) but we should monitor it constantly. To be on the safe side we decided to bring all wire frames inside the clean room to prevent the absorption of moisture in the Rohacell


- We decided not to use ultrasonic cleaning for the second wire frame. If the vacuuming will work to take out the moisture from the first chamber we may try to use the ultrasonic cleaning for the third chamber, but making sure all the holes are reliably sealed. In fact before the first production chamber, we put a testing frame (old design) in the ultrasonic bath and this frame doesn't show measurable conductivity through the Rohacell; most likely it was sealed better.

- Rigid-flex pressing tool: it's back from the machine shop and today Casey will check the flatness. The shoe was narrowed down to 4.5mm.

- Bill is designing the end gussets that hold the package together. His proposal is to make one piece out of aluminum angle (1x1' 2mm thick). We discussed also the possibility to use Be. Eugene: it will be very expensive (1m diameter) because it's toxic when machined, but it has factor of 4 bigger radiation length. Simon will add the Al ring in the simulation to investigate the effects on the calorimeter photon registration efficiency.


- Chris made the new testing card with LEDs on the top; still need a cable to use it.

- There was confusion when populating the PCBs: the bill of materials was in contradiction with the schematics. We populated the boards according the the bill, but it turned out the schematics is the correct one. The HV PCBs were corrected (but as it turned out we had bigger problems with the HV), so later we will correct the signal side as well. Kim created a new version of the material list and Fernando will create a traveler with the correct documentation.

Chamber testing

- Caleb started working on the position measurement problem. His plan is to create software (Visual Basic) to track down the problem and also to improve the procedure.

- The position measurements for the second wire frame showed sigma of ~25 microns. Few of the wires exceeded 100 even 200 microns offsets but were corrected by re-taping the wires. This is the first wire frame after the pin rail was fixed: now it doesn't show the step in the middle but some kink (change in the slope) is visible. Maybe due to the use of the stepper motor. After finishing the second wire frame Bill install the new encoder and hopefully this will fix the problem.